# The Eye as a Window to Systemic Health: A Survey of Retinal Imaging from Classical Techniques to Oculomics

### v1

@@ IV-B2 Diabetes and Diabetic Retinopathy
Diabetes, or in other words, Diabetes Mellitus (DM), is one of the most pressing challenges in the socio-health scenario of the 21st century. It has experienced a dramatic rise in the incidence rate for the past three decades. The international diabetes Federation estimates that, to the tune of 536.5 million, the adult age group suffered from diabetes in the year 2021. This number could reach 738.2 million by 2045 [142].
@@
This disease association is characterised by hypercalcemia, which further exacerbates other complications like CVDS, kidney failure, and vision loss and imposes substantial burdens on the economy and healthcare system [94]. Microvasculature complications damage the small blood vessels linked to the different organs and diseases like Diabetic Nephropathy, Neuropathy and Retinopathy [143]. Microvasculature plays an important role in maintaining tumorigenesis and organ health, and the human eye, i.e., the retinal fundus, serves as a window for in vivo assessment [50]. OCT and OCTA have facilitated the non-invasive assessment of retinal microvasculature [144].
@@
Studies have shown that individuals with diabetes exhibit alterations in retinal vascular calibre, including wider venular diameters, which are associated with an increased risk of developing diabetes and its complications. A comprehensive analysis with 18,771 participants indicated that a larger retinal venular diameter is linked to a higher risk of developing Type-2 diabetes over a median follow-up duration of 10 years, even when controlling for possible confounding factors [145].
@@
Moreover, retinal texture analysis has emerged as a potential tool for the early detection of diabetes-related retinal changes. A recent study utilising OCT-derived texture characteristics found that changes in retinal texture are concurrent with biological retinal changes, suggesting that texture analysis could serve as a biomarker for the early diagnosis of DR [142]. Circulating biomarkers have also been found, which also play a role in predicting diabetic complications. Higher levels of immunocomplexes and immunoglobulin M anti-cardiolipin antibodies have been observed in diabetes patients with vascular complications, suggesting their potential as biomarkers for DR [146].
@@
Similarly, the progression of DR is closely linked to systemic metabolic control, evidenced by some recent studies, which underline the implication of advanced glycation end products (AGEs) and oxidative stress in the development of DR. The accumulation of AGEs in the retinal vasculature triggers the complication of inflammation and endothelial dysfunction; by the same oxidative stress, it worsens the cellular damage [96, 147]. These complications, on the one side, contribute to the progression of DR and, on the other side, reflect the systemic disease related to diabetes, like neuropathy and nephropathy. For example, the presence of DR has been associated with a higher risk of diabetic kidney disease, underscoring the interconnectedness of microvascular complications [148, 149].
@@ IV-B3 Neurodegenerative Diseases
The retina, an extension of the CNS, bridges the functional and structural relationship with the brain and acts as a window to identify early neurodegenerative diseases (NDDs). This unique role of the eye is increasingly considered as a window into the brain, manifesting different NDDs like multiple sclerosis, Parkinson’s, and Alzheimer’s diseases (AD) [3].
Retinal imaging has revealed structural and functional alterations in patients with NDDs.  For example, RNFL thinning and GC layer loss have been observed in individuals with AD. Retinal microvascular analysis provides valuable insights into AD. Studies have shown that individuals with a high genetic risk of AD exhibit reduced macular thickness even before symptoms appear [150], and patients with AD have significantly thinner central macula compared to healthy controls [151]. Additionally, retinal microvascular changes, such as narrower venules and increased vessel tortuosity, may reflect vascular abnormalities associated with AD-related brain pathology, suggesting that retinal imaging could serve as a robust and reliable biomarker for the early detection and monitoring of AD and dementia [152, 153, 154].
Similarly, in Parkinson’s disease, retinal thinning and dopaminergic cell loss in the retina have been linked to disease severity and progression [155, 156]. Currently, OCT has been utilized to quantify these changes, reflecting the progression of neurodegenerative processes impacting the brain. In addition, it has shown notable progress in detecting early signs of multiple sclerosis, with thinning of RNFL linked to disease activity and disability [157] Using imaging techniques other than the retinal and OCT, such as hyperspectral imaging, has further improved the ability to detect subtle retinal changes associated with neurodegeneration. Hyperspectral imaging allows us to visualise the cellular and molecular changes that occur in the retina, unlocking insight into the intricate mechanisms of disease [158]. These recent and other upcoming advances can open the door to a deeper understanding of how these conditions develop and progress, sparking a sense of curiosity and participation among researchers and clinicians.
@@ V Future Directions in Oculomics
    This cannot be covered by the existing public dataset, such as those used in [66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76]. Expanding epidemiological studies, such as cohort and longitudinal approaches, are essential and will cover the genetic sequencing data from the same patients, allowing for cross-comparison of gene expression, imaging biomarkers, and clinical factors. Tracking the same patients over time will help us to understand the underlying mechanism and result in a reliable outcome.
    Likewise, this will create retinal imaging datasets that include multi-omics data, such as genomics, transcriptomics, proteomics, and metabolomics—which are crucial to uncovering disease mechanisms at a systems level. This approach will allow researchers to study the genetic inclination to retinal diseases, track disease progression through molecular signatures, and establish links between ocular and systemic health. This chain will be interlinked with fostering interdisciplinary collaboration in areas like ophthalmology, data science, computer science, bioinformatics, molecular biology, and life sciences, opening a door for the success of oculomics. Similarly, the easily accessible multi-dataset (open access) will be a gold mine for institutional and healthcare research. The policymakers engage to ensure its practical implementation regarding data privacy and security to innovate the advancement in oculomics.
@@
    Expanding the retinal imaging capabilities to work with a diverse range of imaging, like OCT, OCTA, OA, fundus and infrared imaging. The Automorph package still has the same limitation: they are only suitable for the fundus images and cannot be relied on if its modalities are other than this. A diverse range of images should be integrated, which will be suitable for all the modalities of retinal images. This improvement will help us with the subtle retinal changes in the cellular and vasculature network of the eyes and act as ADS for localised and systemic diseases like hypertension, diabetes and neurodegenerative conditions.
@@
    Foundation model and domain shifting are the future of medical image analysis; these two are scarce in oculomics. Domain shifting is hard to correlate with the FP as it is well suitable for the OCT, and this can be achieved by shifting histoimaging of different tissues into the domain of eye tissue. However, the question arises as to why this model cannot be applied to all the modalities of optic systems. However, there is promise in this methodology, and it could be a popular approach in the future. Subsequently, the foundation model is trained on the enormous, diverse dataset and, by the same fine-tuning through a large number of tasks, produces reliable models. This model is found rare in the study of the retina and can be quite helpful in promoting ophthalmology. For example, the retina is linked with different diseases, including CVD. Finding the biomarkers for CVD with respect to retinal imaging. Creating a different model outcome like the features of carotid arteries correlates with the retinal imaging, the same for the serum lipidomics, sleep monitoring and oral microbiome dataset. Exploring the dataset directly connected with CVD and is associated with retinal imaging through the statistical model. Once a linkage is found and at the end, we add these models via the foundation model. This mechanism will be another way of contributing to eye research. In addition, this will maximise the transition of the research setting to routine clinical usage and could be expandable to portable retinal imaging devices to work as an ADS tool, which will also be helpful for the easily accessible remote regions of third-world countries.
@@
    LLM is another buzzword in today’s era and is trained on almost the world’s puzzle. However, there is still room for oculomics, and we should focus on applications like the NYUTron [159], which is trained on 750,000 patients for routine clinical notes. Expanding to eye research can mitigate the burden on health practitioners. Likewise, the LLM medical specialist model Llava-Med [160] and PathChat [161] demonstrate promising biomedical and histopathology imaging capabilities. There is a dire need for oculomics research communities to design an application specific to retinal systems to uncover complex patterns in one go.

### v2

@@ IV-B2 Diabetes and Diabetic Retinopathy
Diabetes, or in other words, Diabetes Mellitus, is one of the most pressing challenges in the socio-health scenario of the 21st century. It has experienced a dramatic rise in the incidence rate for the past three decades. The international diabetes Federation estimates that, to the tune of 536.5 million, the adult age group suffered from diabetes in the year 2021. This number could reach 738.2 million by 2045 Sun et al. (2022).
@@
This disease association is characterised by hypercalcemia, which further exacerbates other complications like CVDS, kidney failure, and vision loss and imposes substantial burdens on the economy and healthcare system Azeem et al. (2022). Microvasculature complications damage the small blood vessels linked to the different organs and diseases like Diabetic Nephropathy, Neuropathy and Retinopathy Bereda (2022). Microvasculature plays an important role in maintaining tumorigenesis and organ health, and the human eye, i.e., the retinal fundus, serves as a window for in vivo assessment Zekavat et al. (2022). OCT and OCTA have facilitated the non-invasive assessment of the retinal microvasculature Oliveira et al. (2025).
@@
Studies have shown that individuals with diabetes exhibit alterations in retinal vascular caliber, including wider venular diameters, which are associated with an increased risk of developing diabetes and its complications. A comprehensive analysis with 18,771 participants indicated that a larger retinal venular diameter is linked to a higher risk of developing Type-2 diabetes over a median follow-up duration of 10 years, even when controlling for possible confounding factors Sabanayagam et al. (2015).
@@
Moreover, retinal texture analysis has emerged as a potential tool for the early detection of diabetes-related retinal changes. A recent study utilising OCT-derived texture characteristics found that changes in retinal texture are concurrent with biological retinal changes, suggesting that texture analysis could serve as a biomarker for the early diagnosis of DR Sun et al. (2022). Circulating biomarkers have also been found, which also play a role in predicting diabetic complications. Higher levels of immunocomplexes and immunoglobulin M anti-cardiolipin antibodies have been observed in diabetes patients with vascular complications, suggesting their potential as biomarkers for DR Simó-Servat et al. (2016).
@@
Similarly, the progression of DR is closely linked to systemic metabolic control, evidenced by some recent studies, which underline the implication of advanced glycation end products (AGEs) and oxidative stress in the development of DR. The accumulation of AGEs in the retinal vasculature triggers the complication of inflammation and endothelial dysfunction; by the same oxidative stress, it worsens the cellular damage Wei et al. (2022); Zhan (2023). These complications, on the one side, contribute to the progression of DR and, on the other side, reflect the systemic disease related to diabetes, like neuropathy and nephropathy. For example, the presence of DR has been associated with a higher risk of diabetic kidney disease, underscoring the interconnectedness of microvascular complications Yamanouchi et al. (2019); Fang et al. (2023).
@@ IV-B3 Neurodegenerative Diseases
The retina, an extension of the central nervous system, offers a unique opportunity to visualize brain health in vivo. Acting as a “window to the brain,” retinal imaging has revealed structural and functional alterations that mirror pathological processes in Alzheimer’s disease (AD), Parkinson’s disease (PD), multiple sclerosis (MS), and even cerebrovascular disease. This dual role, as both a mirror and a biomarker, has increasingly positioned oculomics as a powerful approach to detect early neurodegenerative changes before they manifest clinically Honavar (2022).
In Alzheimer’s disease, thinning of the retinal nerve fiber layer and the loss of the ganglion cell layer (GCL) are among the most consistent findings. Studies have shown that individuals at high genetic risk for AD exhibit reduced macular thickness even before cognitive symptoms emerge López-Cuenca et al. (2020). Patients with AD also present with significantly thinner central macula compared with healthy controls Farzinvash et al. (2022). Retinal microvascular abnormalities, such as narrower venules and increased tortuosity, further suggest a vascular contribution to AD-related pathology Ong et al. (2014); Cheung et al. (2014); Czakó et al. (2020). AI-based models analyzing retinal features have achieved promising results: one deep learning study using over 12,949 fundus images from 648 AD patients achieved diagnostic accuracies ranging from 0.80 to 0.92 Cheung et al. (2022). However, while these biomarkers are strongly correlative, their independent predictive value beyond established markers such as amyloid PET or cerebrospinal fluid assays remains to be validated.
In Parkinson’s disease, retinal changes also mirror underlying brain pathology. OCT studies have revealed retinal thinning and dopaminergic cell loss, both correlating with disease severity and progression Satue et al. (2014); Elanwar et al. (2023). Beyond these structural alterations, functional markers are also being explored. Eye-tracking studies have identified prolonged saccadic latency and reduced response accuracy in PD patients, with machine learning models achieving AUC values of 0.73–0.93 in distinguishing patients from healthy controls Miles et al. (2024); Maleki et al. (2024). Although these findings remain largely correlative, they highlight the potential of combining structural retinal imaging with functional eye-movement analysis to improve early, non-invasive monitoring of neurodegenerative disease. In Alzheimer’s disease, deep learning applied to eye-tracking data has likewise shown that visual attention heatmaps can distinguish patients from controls with performance metrics of 0.84 accuracy and 0.90 AUC Zuo et al. (2024). This provides early evidence that non-invasive eye-movement behaviours could serve as functional biomarkers for cognitive decline.
In multiple sclerosis, evidence is stronger and more validated: Longitudinal OCT studies consistently demonstrate thinning of the retinal RNFL and GCL as biomarkers of disease activity and disability progression Bsteh et al. (2024). These structural markers have been further supported by machine learning approaches: support vector machines, ensemble classifiers, and recurrent neural networks applied to OCT parameters have achieved AUC values ranging from 0.82 to 0.88 for MS diagnosis and prediction of progression Montolío et al. (2021). More recently, Lopez-Dorado et al. (2021) demonstrated that CNNs trained in OCT thickness maps of sweep source, augmented with synthetic data using a deep generative adversarial network, could distinguish newly diagnosed MS patients from controls with near-perfect sensitivity and specificity (both 1.0). The retinal structures contributing most to discrimination included the ganglion cell layer plus inner plexiform layer (GCL+), the extended ganglion cell complex including RNFL and inner plexiform layer (GCL++), and the complete retina. These findings establish retinal imaging as both a structural and a computational biomarker platform to monitor MS, complementing its role in other neurodegenerative disorders.
Finally, cerebrovascular disease links further underscore the eye–brain connection. Longitudinal population-based studies consistently show that retinal vascular traits are predictive of stroke risk. In the Rotterdam Study, wider venular caliber was independently associated with both cerebral infarction and intracerebral hemorrhage, with hazard ratios of 1.28 (95% CI: 1.13–1.46) and 1.53 (95% CI: 1.09–2.15), respectively, while associations with narrower arteriolar caliber were weaker and borderline significant Wieberdink et al. (2010). The Multi-Ethnic Study of Atherosclerosis (MESA) extended these findings to a diverse U.S. cohort, demonstrating that narrower arteriolar caliber (HR = 3.01, 95% CI: 1.29–6.99) and retinopathy in non-diabetic individuals (HR = 3.07, 95% CI: 1.17–8.09) were independently associated with incident stroke, even after adjustment for atherosclerotic markers such as carotid intima-media thickness and coronary calcium Kawasaki et al. (2012)(Klein et al., 2010). In Asian populations, the Singapore Malay Eye Study showed that larger venular caliber (HR = 3.28, 95% CI: 1.30–8.26) and retinopathy signs predicted stroke events, and that the inclusion of retinal microvascular measures improved risk prediction and reclassification beyond established factors Cheung et al. (2013). The decrease in retinal fractal dimension was similarly associated with stroke risk, with an odds ratio of 1.80 for the venular and 2.28 for the arteriolar networks Wu et al. (2017). More recently, large-scale AI studies have strengthened these associations. A UK Biobank study of 45,161 participants reported that each standard-deviation change in vascular traits, including vessel density, caliber, and branching complexity, was associated with a 9.8–19.0% change in stroke risk. Incorporating these measures into conventional models improved predictive accuracy, increasing the area under the ROC from 0.739 to 0.752 Yusufu et al. (2025). Similarly, the DeepRETStroke framework, trained on hundreds of thousands of retinal photographs across multiple countries, predicted incident stroke with an AUC of 0.901 by detecting silent brain infarctions, outperforming conventional predictors across diverse populations Jiang et al. (2025).
These validated findings illustrate how subtle retinal vascular changes can forecast cerebrovascular accidents years in advance, making them one of the most clinically actionable domains of oculomics. Collectively, these examples show how the retina serves as a biological bridge between the eye and brain. While many associations remain exploratory, validated pathways in multiple sclerosis and stroke demonstrate the clinical promise of oculomics. Ongoing advances in AI, multimodal imaging, and wearable eye-tracking technologies are expected to strengthen this bridge, enabling earlier detection, more accurate risk stratification, and deeper insights into the mechanisms of neurodegenerative disease.
@@ V Future Directions in Oculomics
    This cannot be covered by the existing public dataset, such as those used in Hoover et al. (2000); Staal et al. (2004); Kauppi et al. (2007); Owen et al. (2009); Budai et al. (2013); Decenciere et al. (2013); Decencière et al. (2014); Dugas et al. (2015); Porwal et al. (2018); Karthik et al. (2019); Li et al. (2019). Expanding epidemiological studies, such as cohort and longitudinal approaches, are essential and will cover the genetic sequencing data from the same patients, allowing for cross-comparison of gene expression, imaging biomarkers, and clinical factors. Tracking the same patients over time will help us to understand the underlying mechanism and result in a reliable outcome.
    Likewise, this will create retinal imaging datasets that include multi-omics data, such as genomics, transcriptomics, proteomics, and metabolomics, which are crucial to uncovering disease mechanisms at a systems level. This approach will allow researchers to study the genetic inclination to retinal diseases, track disease progression through molecular signatures, and establish links between ocular and systemic health. This chain will be interlinked with fostering interdisciplinary collaboration in areas like ophthalmology, data science, computer science, bioinformatics, molecular biology, and life sciences, opening a door for the success of oculomics. Similarly, the easily accessible multi-dataset (open access) will be a gold mine for institutional and healthcare research. The policymakers engage to ensure the practical implementation of data privacy and security to innovate the advancement in oculomics.
@@
    Expanding the retinal imaging capabilities to work with a diverse range of imaging, namely OCT, OCTA, AO-OCT, fundus and infrared imaging. The Automorph package still has the same limitation: they are only suitable for the fundus images and cannot be relied on if its modalities are other than this. A diverse range of images should be integrated, which will be suitable for all the modalities of retinal images. This improvement will help us with the subtle retinal changes in the cellular and vasculature network of the eyes and act as ADS for localised and systemic diseases like hypertension, diabetes and neurodegenerative conditions.
@@
    Foundation model and domain shifting are the future of medical image analysis; these two are scarce in oculomics. Domain shifting is hard to correlate with the CFP as it is well suitable for the OCT, and this can be achieved by shifting histoimaging of different tissues into the domain of eye tissue. However, the question arises as to why this model cannot be applied to all the modalities of optic systems. However, there is promise in this methodology, and it could be a popular approach in the future. Subsequently, the foundation model is trained on the enormous, diverse dataset and, by the same fine-tuning through a large number of tasks, produces reliable models. This model is found rare in the study of the retina and can be quite helpful in promoting ophthalmology. For example, the retina is linked with different diseases, including CVD. Finding the biomarkers for CVD with respect to retinal imaging. Creating a different model outcome like the features of carotid arteries correlates with the retinal imaging, the same for the serum lipidomics, sleep monitoring and oral microbiome dataset. Exploring the dataset directly connected with CVD and is associated with retinal imaging through the statistical model. Once a linkage is found and at the end, we add these models via the foundation model. This mechanism will be another way of contributing to eye research. In addition, this will maximise the transition of the research setting to routine clinical usage and could be expandable to portable retinal imaging devices to work as an ADS tool, which will also be helpful for the easily accessible remote regions of third-world countries.
@@
    LLM is another buzzword in today’s era and is trained on almost the world’s puzzle. However, there is still room for oculomics, and we should focus on applications like the NYUTron Jiang et al. (2023), which is trained on 750,000 patients for routine clinical notes. Expanding to eye research can mitigate the burden on health practitioners. Likewise, the LLM medical specialist model Llava-Med Li, Wong, Zhang, Usuyama, Liu, Yang, Naumann, Poon & Gao (2024) and PathChat Lu et al. (2023) demonstrate promising biomedical and histopathology imaging capabilities. There is a dire need for oculomics research communities to design an application specific to retinal systems to uncover complex patterns in one go.